OpenStudy (anonymous):

What is the geometric mean of 3 and 21? 63 31.5 3 square root 7 7 square root 3

Parth (parthkohli):

\(\Rightarrow G.M(3,21) = \sqrt{3 \times 21} = \sqrt{63}\)

OpenStudy (anonymous):

WAIT isnt it the third one

Parth (parthkohli):

Which can be simplified as \(\sqrt{7 \times 3 \times 3} = 3\sqrt{7}\)

Parth (parthkohli):

Yes it is :)

OpenStudy (anonymous):

:)

OpenStudy (ganpat):

Geometric mean is given by = \[\sqrt{(3 * 21)}\]. = sq rt of 63 = 3 \[\sqrt{7}\] = 7.93